If the temperature of a 1.5 L reaction vessel containing Argon gas is decreased from 5.60°C to -12.9°C and the pressure remains constant, what is the new volume of gas?
1
Expert's answer
2020-06-22T09:16:49-0400

Assuming pressure is constant:

"\frac{V1}{T1}=\frac{V2}{T2}"


T1=5,60 ° С =278.75K

T2=12,9 ° С=260.25K


"\frac{1.5}{278.75}=\frac{V2}{260.25}"


V2=1.4 L